OpenStudy (anonymous):

answer check plse: factor completely c^2 + 81 - 18c = (c-9)^2 ans

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[c^2-18c+81=(c-9)^2\]

OpenStudy (anonymous):

checking answes factor the trinomial b^2 - 7b + 10 = (b-5)(b-2)ans